OpenStudy (anonymous):

What is the maximum current delivered to a circuit containing a4.20 µF capacitor when it isconnected across the following outlets? (a) a North American outlet havingΔVrms = 120 V, f = 60.0 Hz

OpenStudy (radar):

Calculate the capacitive reactance at 60 Hz\[1 \over2\pi \times60.0\times4.2\times10^{-6}\]=631.567 Ohms of capacitance reactance. Maximum current =120/631.567=0.19 amp